# Sample Compression Scheme Reductions

### v2

@@ 1.1 Our Results
  • We construct an $\epsilon$-approximate sample compression scheme for real-valued concept classes of size
    $O\mathopen{}\left(f(d_{\mathrm{P}}(\mathcal{C}))\log\frac{1}{\epsilon}\right)$ for the $\ell_{\infty}$ loss, and $O\mathopen{}\left(f(d_{\mathrm{P}}(\mathcal{C}))\frac{1}{p}\log\frac{1}{\epsilon}\right)$ for the $\ell_{p}$ loss, $p\in[1,\infty)$ (Theorem 4.2).
@@ 3 Compression for Multiclass Classification
In this section, we tackle the problem of multiclass compression by reducing it to the binary setting. In multiclass classification, the finiteness of the graph dimension of a concept class characterizes learnability when the label set is finite (albeit with nonoptimal sample complexity in general). David et al. [2016] demonstrated that, in such cases, a sample compression scheme of size $2^{O\mathopen{}\left(d_{\mathrm{G}}\right)}$ can be constructed, with the compression size notably independent of the sample size. However, Pabbaraju [2024] showed that when the label set is infinite, there exist concept classes where any sample compression scheme must grow at least logarithmically with the sample size.
@@
A set of points $x_{1},\ldots,x_{n}\in\mathcal{X}$ is G-shattered by $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$ if there exist $y_{1},\ldots,y_{n}\in\mathcal{Y}$ such that
@@
where $g_{c}:\mathcal{X}\times\mathcal{Y}\rightarrow\{0,1\}$ is defined such that $g_{c}(x,y)=\mathbbm{1}[c(x)=y]$. In the following, we construct a multiclass sample compression scheme for classes of finite labels and finite graph dimension, via a reduction to the binary setting.
@@
Suppose
that for binary concept classes with finite VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for multiclass concept classes with a finite label set $|\mathcal{Y}|$ and a graph dimension $d_{\mathrm{G}}<\infty$, there exists a sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\log|\mathcal{Y}|\right)$.
@@
Suppose $\mathcal{C}_{\mathcal{Y}}$ has a binary compression scheme ($\kappa_{b}$,$\rho_{b}$), then we construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given a dataset $S=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ realizable by $\mathcal{C}$, construct $\kappa(S)$ as follows. Inflate $S$ to $\mathcal{S_{\mathcal{Y}}}$. Note that $S_{\mathcal{Y}}$ is realizable by $\mathcal{C}_{\mathcal{Y}}$, and $d_{\mathrm{VC}}=d_{\mathrm{G}}$. We can apply $\kappa_{b}$ to $S_{\mathcal{Y}}$ to get a compression of size $f(d_{\mathrm{G}})$. The compression points will be of the form $((x,y),z),z\in\{0,1\}$. For the points where $z=1$, we have that $(x,z)=(x_{i},y_{i})$ for some $i$, so we can add that to $\kappa(S)$, contributing $1$ for each. For the points where $z=0$, we need $\log|\mathcal{Y}|$ bits to add that to $\kappa(S)$. Thus, our compression size will be $\leq f(d_{\mathrm{G}})+f(d_{\mathrm{G}})\log|\mathcal{Y}|=O(f(d_{\mathrm{G}})\log|\mathcal{Y}|)$. Reconstruction: Our compression has enough information for us to retrieve the result of $\kappa_{b}(\mathcal{C}_{\mathcal{Y}})$. We can directly apply $\rho_{b}$ on this to get the desired result.
For classes with graph dimension $1$ we can get a tighter result, but for general concept classes and binary compression schemes it is an open problem whether we can remove $\log\mathopen{}\left|\mathcal{Y}\right|$ from the compression in Theorem 3.2.
@@ 3.1 Additional Assumption: Existence of Proper or Majority Vote Binary Compression
A binary function $f:\mathcal{X}\rightarrow\{0,1\}$ is a majority of concepts from $\mathcal{C}\subseteq\mathopen{}\left\{0,1\right\}^{\mathcal{X}}$ if there is a $\mathcal{C}_{f}\subseteq\mathcal{C}$ such that for all $x\in\mathcal{X}$, $f(x)=\mathrm{Maj}(c(x):c\in\mathcal{C}_{f})$ where $\mathrm{Maj}(\cdot)$ takes in a sequence and returns the majority element (picking zero to break ties). A compression scheme is a majority vote compression scheme if for every finite $S\in(\mathcal{X}\times\mathcal{Y})^{*}$, $\rho(\kappa(S))$ outputs a majority of concepts from $\mathcal{C}$.
@@
Suppose any binary concept class $\mathcal{C}$ with VC dimension $d_{\mathrm{VC}}<\infty$ has a compression scheme of size $f(d_{\mathrm{VC}})$ which is a proper compression scheme or a majority vote compression scheme (see Definition 3.4). Then any multiclass concept class (label set is allowed to be of infinite size) with a finite graph dimension $d_{\mathrm{G}}<\infty$ admits a compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$.
In particular, we recover the best known bounded-size multiclass compression scheme of size $2^{O\mathopen{}\left(d_{\mathrm{G}}\right)}$ [David et al., 2016], via a reduction to binary compression scheme.
@@
Suppose $\mathcal{C}_{\mathcal{Y}}$ (Equation 2) has a binary compression scheme ($\kappa_{b}$,$\rho_{b}$), which is either a proper compression scheme or a majority vote compression scheme. Consider $g=\rho_{b}(T)$ for any $T$ in the image of $\kappa_{b}$ and any possible finite realizable $S$ given to $\kappa_{b}$ as an input. We claim that for all $x\in\mathcal{X}$, $g(x,y)$ equals $1$ for at most one $y$. To show this, we consider two cases: 1) $g$ is proper, and 2) $g$ is a majority of concepts from the concept class. In the first case, since $g$ is proper in $\mathcal{C}_{\mathcal{Y}}$, $g(x,y)$ must equal $1$ for exactly one $y\in\mathcal{Y}$. Now consider the second case, where $g$ is a majority of concepts $\mathcal{C}_{g}\subseteq\mathcal{C}$. Notice that for any fixed $x\in\mathcal{X}$, $c(x,y)=1$ for exactly $|\mathcal{C}_{g}|$ pairs $(c,y)$ in $\mathcal{C}_{g}\times\mathcal{Y}$ (since for any $c\in\mathcal{C}_{g}$, there exists exactly one $y\in\mathcal{Y}$ such that $c(x,y)=1$). Thus, in order for $g(x,y)$ to be $1$, $c(x,y)$ must be $1$ for strictly more than $|\mathcal{C}_{g}|/2$, $c\in\mathcal{C}_{g}$. Thus, $c(x,y)=1$ for at most one $y\in\mathcal{Y}$.
We now construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given realizable $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, let $T=\{((x_{i},y_{i}),1):i\in[n]\}$. We let our compression $\kappa(S)$ return the $(x,y)$ pairs from $\kappa_{b}(T)$, which have size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$. Reconstruction: We have that $\kappa_{b}(T)=\{((x,y),1):(x,y)\in\kappa(S)\}$, so we can immediately recover $\kappa_{b}(T)$. We know that $\rho_{b}(\kappa_{b}(T))$ is correct on all possible $T$. Additionally, since we predict $1$ for at most one $y$ for each $x$, we can conclude that $\rho_{b}(\kappa_{b}(T))$ is correct on all of $S_{\mathcal{Y}}$, as desired. Note that if the reconstruction $\rho_{b}$ is proper, then it will output exactly one $1$ for each $x$. Additionally, it the reconstruction is a majority of learners from the class, where we break ties to favor predicting 00, then it will output at most one $1$ for each $x$.
@@ 3.2 Additional Assumption: Existence of Stable Binary Compression
Sample compression scheme is stable if for any sequence $S$ over $(\mathcal{X}\times\mathcal{Y})$, and any $T:\kappa(S)\subseteq T\subseteq S$, it holds that $\kappa(T)=\kappa(S)$  [^2] Some notations: If there is a $\subseteq$ sign and any of the sides is of the form $(S,b)$ for a sequence $S$ and a bitstring $b$, we can interpret it to be $S$. For example, $(S,b)\subseteq T$ will be interpreted as $S\subseteq T$. Also, we will sometimes apply $\kappa$ to something of the form $(T,b)$ (for example, something like $\kappa(\kappa(S))$). In this case, we can interpret it to be $\kappa(T)$. .
@@
Suppose that for binary concept classes with finite VC dimension $d_{\mathrm{VC}}<\infty$, there exists a stable sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for multiclass concept classes with a finite graph dimension $d_{\mathrm{G}}<\infty$ and finite label space, there exists a stable sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$.
@@
$\kappa^{\prime}_{b}(S_{\mathcal{Y}}):=\{((x_{i},y),\mathbbm{1}[y=y_{i}]):\exists z,w:((x_{i},z),w)\in\kappa_{b}(S_{\mathcal{Y}}),y\in\mathcal{Y}]\},$
i.e., it is an inflated $\kappa_{b}(S)$ to include all the labels for each $x$. Since the compression scheme is stable, and since $\kappa_{b}(S_{\mathcal{Y}})\subseteq\kappa_{b}^{\prime}(S_{\mathcal{Y}})\subseteq S_{\mathcal{Y}}$, we must have that $\kappa_{b}(\kappa_{b}^{\prime}(S_{\mathcal{Y}}))=\kappa_{b}(S_{\mathcal{Y}})$. We can let $\kappa(S)$ return the $(x_{i},y_{i})$ pairs such that $x_{i}$ is a member of $\kappa_{b}(S_{\mathcal{Y}})$. This will have size $O(f(d_{\mathrm{G}}))$. Reconstruction: Given $\kappa(S)$, we can reconstruct $\kappa^{\prime}_{b}(S_{\mathcal{Y}})$ by inflating the dataset to include all labels for each $x$. Then, we can apply $\kappa_{b}$ to get $\kappa_{b}(S_{\mathcal{Y}})$ (using the fact that it is a stable compression scheme), and then apply $\rho_{b}$ to get the desired compression scheme.
@@
Theorem 3.9 (Multiclass, Infinite Labels, Reductions with Stable Infinitized Compression Schemes)
@@
Given a class on $\{0,1\}^{(\mathcal{X}\times\mathcal{Y})}$, define an inflated compression scheme by the following two functions:
@@
Additionally, it must hold that for any realizable sequence $S\in(\mathcal{X}\times\mathcal{Y}\times\{0,1\})^{\infty}$ where
$\{x\in\mathcal{X}:\exists y,z\ s.t.((x,y),z)\in S\}$ is finite, $\rho(\kappa(S))(x)=y\mathrm{\text{ for all }}(x,y)\in S$. The inflated sample compression scheme is of size $k$ if for any realizable sequence $S\in(\mathcal{X}\times\mathcal{Y}\times\{0,1\})^{\infty}$ where $\{x\in\mathcal{X}:\exists y,z\ s.t.((x,y),z)\in S\}$ is finite, for $\kappa(S)=(S^{\prime},b)$, $|S^{\prime}|+|b|\leq k$.
Theorem 3.11 (Multiclass, Infinite Labels, Reductions with Stable Inflated Compression Schemes)
@@
Example 3.12 (A Concept Class with Inflated Stable Compression Scheme: Piecewise Thresholds)

### v3

@@ Abstract
Keywords: Sample Compression Schemes, PAC Learning, Binary Classification, Multiclass Classification, Regression, Adversarially Robust Learning.
@@ 1.1 Our Results
  • We construct an $\epsilon$-approximate sample compression scheme for real-valued concept classes of size $O\mathopen{}\left(f(d_{\mathrm{P}}(\mathcal{C}))\log\frac{1}{\epsilon}\right)$ for the $\ell_{\infty}$ loss, and $O\mathopen{}\left(f(d_{\mathrm{P}}(\mathcal{C}))\frac{1}{p}\log\frac{1}{\epsilon}\right)$ for the $\ell_{p}$ loss, $p\in[1,\infty)$ (Theorem 4.2).
@@ 3 Compression for Multiclass Classification
In this section, we tackle the problem of multiclass compression by reducing it to the binary setting. In multiclass classification, the finiteness of the graph dimension of a concept class characterizes learnability when the label set is finite (albeit with non-optimal sample complexity in general). David et al. [2016] demonstrated that, in such cases, a sample compression scheme of size $2^{O\mathopen{}\left(d_{\mathrm{G}}\right)}$ can be constructed, with the compression size notably independent of the sample size. However, Pabbaraju [2024] showed that when the label set is infinite, there exist concept classes where any sample compression scheme must grow at least logarithmically with the sample size.
@@
A set of points
$x_{1},\ldots,x_{n}\in\mathcal{X}$ is G-shattered by $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$ if there exist $y_{1},\ldots,y_{n}\in\mathcal{Y}$ such that
@@
where $g_{c}:\mathcal{X}\times\mathcal{Y}\rightarrow\{0,1\}$ is defined such that $g_{c}(x,y)=\mathbbm{1}[c(x)=y]$. This inflation operation transforms the multiclass prediction problem into a binary classification problem: for each original example $(x_{i},y_{i})$, we create $|\mathcal{Y}|$ binary-labeled examples, where the positive label indicates the correct class. Similarly, each multiclass concept $c$ is transformed into a binary concept $g_{c}$ that identifies correct label predictions. In the following, we construct a multiclass sample compression scheme for classes of finite labels and finite graph dimension, via a reduction to the binary setting.
@@
Suppose that for binary concept classes with finite VC dimension $d_{\mathrm{VC}}<\infty$, there exists a sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for multiclass concept classes with a finite label set $|\mathcal{Y}|$ and a graph dimension $d_{\mathrm{G}}<\infty$, there exists a sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\log|\mathcal{Y}|\right)$.
@@
Suppose $\mathcal{C}_{\mathcal{Y}}$ has a binary compression scheme ($\kappa_{b}$,$\rho_{b}$), then we construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given a dataset $S=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ realizable by $\mathcal{C}$, construct $\kappa(S)$ as follows. Inflate $S$ to $\mathcal{S_{\mathcal{Y}}}$. Note that $S_{\mathcal{Y}}$ is realizable by $\mathcal{C}_{\mathcal{Y}}$, and $d_{\mathrm{VC}}=d_{\mathrm{G}}$. We can apply $\kappa_{b}$ to $S_{\mathcal{Y}}$ to get a compression of size $f(d_{\mathrm{G}})$. The compression points will be of the form $((x,y),z),z\in\{0,1\}$. For the points where $z=1$, we have that $(x,y)=(x_{i},y_{i})$ for some $i$, so we can add that to $\kappa(S)$, contributing $1$ for each. For the points where $z=0$, we need $\log|\mathcal{Y}|$ bits to add that to $\kappa(S)$. Thus, our compression size will be $\leq f(d_{\mathrm{G}})+f(d_{\mathrm{G}})\log|\mathcal{Y}|=O(f(d_{\mathrm{G}})\log|\mathcal{Y}|)$. Reconstruction: Our compression has enough information for us to retrieve the result of $\kappa_{b}(\mathcal{C}_{\mathcal{Y}})$. We can directly apply $\rho_{b}$ on this to get the desired result.
For classes with graph dimension $1$ we can get a tighter result, but for general concept classes and binary compression schemes it is an open problem whether we can remove the $\log\mathopen{}\left|\mathcal{Y}\right|$ factor from the compression size in Theorem 3.2.
@@ 3.1 Additional Assumption: Existence of Proper or Majority Vote Binary Compression
A binary function $f:\mathcal{X}\rightarrow\{0,1\}$ is a majority of concepts from $\mathcal{C}\subseteq\mathopen{}\left\{0,1\right\}^{\mathcal{X}}$ if there exists a finite $\mathcal{C}_{f}\subseteq\mathcal{C}$ such that for all $x\in\mathcal{X}$, $f(x)=\mathrm{Maj}(c(x):c\in\mathcal{C}_{f})$, where $\mathrm{Maj}(\cdot)$ takes in a sequence and returns the majority element (picking zero to break ties). A compression scheme is a majority vote compression scheme if for every finite $S\in(\mathcal{X}\times\mathcal{Y})^{*}$, $\rho(\kappa(S))$ outputs a majority of concepts from $\mathcal{C}$.
@@
Suppose any binary concept class $\mathcal{C}$ with VC dimension $d_{\mathrm{VC}}<\infty$ has a compression scheme of size $f(d_{\mathrm{VC}})$ that is either a proper compression scheme or a majority vote compression scheme (see Definition 3.4). Then any multiclass concept class (allowing infinite label sets) with a finite graph dimension $d_{\mathrm{G}}<\infty$ admits a compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$.
In particular, we recover the best known bounded-size multiclass compression scheme of size $2^{O\mathopen{}\left(d_{\mathrm{G}}\right)}$ [David et al., 2016], via a reduction to binary compression.
@@
Suppose $\mathcal{C}_{\mathcal{Y}}$ (Equation 2) has a binary compression scheme ($\kappa_{b}$,$\rho_{b}$), which is either a proper compression scheme or a majority vote compression scheme. Consider $h=\rho_{b}(T)$ for any $T$ in the image of $\kappa_{b}$ and any possible finite realizable $S$ given to $\kappa_{b}$ as an input. We claim that for all $x\in\mathcal{X}$, $h(x,y)$ equals $1$ for at most one $y$. To show this, we consider two cases: 1) $h$ is proper, and 2) $h$ is a majority of concepts from the concept class. In the first case, since $h$ is proper in $\mathcal{C}_{\mathcal{Y}}$, $h(x,y)$ must equal $1$ for exactly one $y\in\mathcal{Y}$. Now consider the second case, where $h$ is a majority of concepts $\mathcal{C}_{h}\subseteq\mathcal{C}$. Notice that for any fixed $x\in\mathcal{X}$, $c(x,y)=1$ for exactly $|\mathcal{C}_{h}|$ pairs $(c,y)$ in $\mathcal{C}_{h}\times\mathcal{Y}$ (since for any $c\in\mathcal{C}_{h}$, there exists exactly one $y\in\mathcal{Y}$ such that $c(x,y)=1$). Thus, in order for $h(x,y)$ to be $1$, $c(x,y)$ must be $1$ for strictly more than $|\mathcal{C}_{h}|/2$ concepts $c$ from $\mathcal{C}_{h}$. Therefore, $h(x,y)=1$ for at most one $y\in\mathcal{Y}$.
We now construct a compression scheme $(\kappa,\rho)$ for $\mathcal{C}$ as follows. Compression: Given realizable $S=(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, let $T=\{((x_{i},y_{i}),1):i\in[n]\}$. We let our compression $\kappa(S)$ return the $(x,y)$ pairs from $\kappa_{b}(T)$, which have size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$. Reconstruction: We have that $\kappa_{b}(T)=\{((x,y),1):(x,y)\in\kappa(S)\}$, so we can immediately recover $\kappa_{b}(T)$. Since $\rho_{b}(\kappa_{b}(T))$ is correct on all possible $T$ and predicts $1$ for at most one $y$ for each $x$, we can conclude that $\rho_{b}(\kappa_{b}(T))$ is correct on all of $S_{\mathcal{Y}}$. Note that if the reconstruction $\rho_{b}$ is proper, then it will output exactly one $1$ for each $x$. Additionally, it the reconstruction is a majority of learners from the class, where we break ties to favor predicting 00, then it will output at most one $1$ for each $x$.
@@ 3.2 Additional Assumption: Existence of Stable Binary Compression
A sample compression
scheme is stable if for any sequence $S$ over $(\mathcal{X}\times\mathcal{Y})$, and any $T:\kappa(S)\subseteq T\subseteq S$, it holds that $\kappa(T)=\kappa(S)$  [^2] Some notations: If there is a $\subseteq$ sign and any of the sides is of the form $(S,b)$ for a sequence $S$ and a bitstring $b$, we can interpret it to be $S$. For example, $(S,b)\subseteq T$ will be interpreted as $S\subseteq T$. Also, we will sometimes apply $\kappa$ to something of the form $(T,b)$ (for example, something like $\kappa(\kappa(S))$). In this case, we can interpret it to be $\kappa(T)$. .
@@
Suppose that for
binary concept classes with finite VC dimension $d_{\mathrm{VC}}<\infty$, there exists a stable sample compression scheme of size $f(d_{\mathrm{VC}})$. Then, for multiclass concept classes with a finite graph dimension $d_{\mathrm{G}}<\infty$ and finite label space, there exists a stable sample compression scheme of size $O\mathopen{}\left(f(d_{\mathrm{G}})\right)$.
@@
$\kappa^{\prime}_{b}(S_{\mathcal{Y}}):=\{((x,y),\mathbbm{1}[y=y_{i}]):x\text{ appears in }\kappa_{b}(S_{\mathcal{Y}}),(x,y_{i})\in S,y\in\mathcal{Y}\},$
i.e., it is an inflated $\kappa_{b}(S_{\mathcal{Y}})$ to include all the labels for each $x$. Since the compression scheme is stable, and since $\kappa_{b}(S_{\mathcal{Y}})\subseteq\kappa^{\prime}_{b}(S_{\mathcal{Y}})\subseteq S_{\mathcal{Y}}$, we must have that $\kappa_{b}(\kappa^{\prime}_{b}(S_{\mathcal{Y}}))=\kappa_{b}(S_{\mathcal{Y}})$. We can let $\kappa(S)$ return the $(x_{i},y_{i})$ pairs such that $x_{i}$ is a member of $\kappa_{b}(S_{\mathcal{Y}})$. This will have size $O(f(d_{\mathrm{G}}))$. Reconstruction: Given $\kappa(S)$, we can reconstruct $\kappa^{\prime}_{b}(S_{\mathcal{Y}})$ by inflating the dataset to include all labels for each $x$ in $\kappa(S)$. Then, we can apply $\kappa_{b}$ to get $\kappa_{b}(S_{\mathcal{Y}})$ (using the fact that it is a stable compression scheme), and then apply $\rho_{b}$ to get the desired compression result.
@@
Theorem 3.9 (Multiclass, $|\mathcal{Y}|=\infty$, Reductions with Stable Infinitized Compression Schemes)
@@
Given a concept class $\mathcal{C}\subset\{0,1\}^{(\mathcal{X}\times\mathcal{Y})}$, define an inflated compression scheme by the following two functions:
@@
Additionally, for any realizable sequence $S\in(\mathcal{X}\times\mathcal{Y}\times\{0,1\})^{\infty}$ where |$|\{x\in\mathcal{X}:\exists y\in\mathcal{Y},z\in\{0,1\}\text{ s.t. }((x,y),z)\in S\}|<\infty$, it must hold that $\rho(\kappa(S))((x,y))=z$ for all $((x,y),z)\in S$. The inflated sample compression scheme is of size $k$ if for any such realizable sequence $S$, where $\kappa(S)=(S^{\prime},b)$, we have $|S^{\prime}|+|b|\leq k$.
Theorem 3.11 (Multiclass, $|\mathcal{Y}|=\infty$, Reductions with Stable Inflated Compression Schemes)
@@
Example 3.12 (Concept Class with Inflated Stable Compression Scheme: Piecewise Thresholds)